# This Version: March 9, 2024

### v1

@@
This Version: March 9, 2024
We present novel monotone comparative statics results for steady state behavior in a dynamic optimization environment with misspecified Bayesian learning. We consider a generalized framework, based on Esponda and Pouzo (2021), wherein a Bayesian learner facing a dynamic optimization problem has a prior on a set of parameterized transition probability functions (models) but is misspecified in the sense that the true process is not within this set. In the steady state, the learner infers the model that best-fits the data generated by their actions, and in turn, their actions are optimally chosen given their inferred model. We characterize conditions on the primitives of the environment, and in particular, over the set of models under which the steady state distribution over states and actions and inferred models exhibit monotonic behavior. Further, we offer a new theorem on the existence of a steady state on the basis of a monotonicity argument. Lastly, we provide an upper bound on the cost of misspecification, again in terms of the primitives of the environment. We demonstrate the utility of our results for several environments of general interest, including forecasting models, dynamic effort-task, and optimal consumption-savings problems.
Journal of Economic Literature Classification Numbers: C1, C5, C6, D8, D9
Key Words: Bayesian learning, Berk-Nash equilibrium, comparative statics, misspecification.
In general, dynamic programs with learning are intractable, i.e., they are not solvable either analytically or numerically, when there is no separability of control and learning. The problem is not only whether a solution exists, but if a solution can be characterized and its properties studied. Koulovitanos, Mirman, and Santugini (2009)
@@ 1 Introduction
Stemming from the foundational contributions of Arrow and Green (1973) and Nyarko (1991), model misspecification in economic environments continues to be of substantive interest for economists. It acknowledges the possibility that economic agents, perhaps due to cognitive limitations or simplified perspectives, may often not know the true model of their complex economic environment and, therefore, learn about it through a set of models that are misspecified in the sense that the true model is not contained in their set of models. A growing body of literature in economics suggests that enriching classical economic environments with misspecified economic agents aligns theoretical predictions closely with the observed behavior. [^1] For example, Farmer, Nakamura, and Steinsson (2023) finds that much of the anomalies concerned with forecasts of professional forecasters can be attributed to them not knowing the true model of the environment.
Dynamic programming offers one robust framework for economists to approach discrete-time economic problems. They are commonly used to analyze instances where economic agents make decisions sequentially in an uncertain environment. The framework applies to a broad spectrum of economic models, from consumption-savings decisions for individuals to investment choices for firms. However, they are known to be intractable, no closed-form solutions, when incorporated with agents who simultaneously make decisions and learn about their environment through their models, where both the decision making and the learning activity mutually affect each other. In this paper, we provide robust predictions on the comparative statics properties of an agent’s behavior within a dynamic programming environment, particularly when they are learning with potentially misspecified models. [^2] As is well-known, comparative statics analysis deals with characterizing conditions for various environments, both static and dynamic, under which decision rules and solution concepts are increasing in the primitives.
@@
$V(s,\mu)=\max_{x\in\mathbb{X}}\Bigl{\{}u\left(s,x\right)+\beta\int_{\mathbb{S}}V\left(s^{\prime},\mu^{\prime}\right){\bar{Q}_{\mu}}\left(ds^{\prime}\mid s,x\right)\Bigl{\}},$ (1)
@@
For a misspecified agent that follows the above setting in choosing their sequence of actions and updating their sequence of beliefs over models, EP predicts its steady-state behavior in terms of a solution concept called the Berk-Nash equilibrium. It is an equilibrium distribution over states and actions corresponding to which the misspecified agent infers the model (or a set of models) that ‘best fits’ the true model. In turn, given the inferred model, the chosen actions are optimal for each state. Furthermore, the equilibrium distribution over states and actions gives rise to a stationary Markov process. The ‘best-fit’ formalization is in the sense of a minimum weighted Kullback-Liebler divergence, with weights determined by the equilibrium distribution over states and actions. [^4] The Kullback-Liebler (KL) divergence measures the difference between two probability distributions in terms of relative entropy; in our case, between the models and the true process; see Cover and Thomas (2005) for details. Given this prediction of the long-run behavior of a misspecified Bayesian agent in terms of a Berk-Nash equilibrium and its corresponding inferred model, we ask the following comparative statics question: how do the equilibrium objects respond to changes in the economic primitives? Phrased differently, what are the requirements on the primitives of the economic environment so that a misspecified agent’s stationary distribution and corresponding best-fit models exhibit monotone comparative statics properties?
@@
This question is of general theoretical interest for several reasons. First, models of dynamic programming incorporated with learning, whether correctly specified or misspecified, such as the one described in Equation (1), are known to be intractable. [^5] With learning incorporated in dynamic programming models, the agent is an active learner and processes incoming data in order to infer the models of their unknown environment. At the same time, they also take decisions. These two functions of the agent are intertwined. In an important paper on learning in stochastic growth models, Koulovatianos, Mirman, and Santugini (2009) (hereafter, KMS) note that the intractability arises primarily for two reasons: (a) the curse of dimensionality problem, which occurs because of incorporation of beliefs about the models in the state space, thus making the state space very large and computationally cumbersome, and (b) with Bayesian updating, the prior and the posterior over models potentially belong to different families, thereby hampering analytical and computational tractability. [^6] As KMS and EP note, the curse of dimensionality does not arise if the beliefs are over a finite parameter set. However, for parameter spaces such as the real line, it is a cause for concern. Also note that, unlike KMS, in our framework, choosing actions also affects the flow of information, thus further complicating matters. By focusing on the long-run prediction for a misspecified dynamic program with a precise characterization of the distribution over states and actions, as well as the limiting posterior (beliefs) over models, we abstract away from the intractability problems and thereby make predictions about their steady-state properties even when their closed-form solutions may not exist. Thus, in this paper, we expand the theory of dynamic programming with misspecified learning, potentially rendering it applicable across a wide spectrum of economic domains.
@@
Second, comparative statics for misspecified learning environments are typically not well-understood in the existing literature on misspecification in economic theory. This traces itself to the issue that with concurrent decision making and learning, there is often no separability between the two, as is illustrated in Equation (1). We tackle this problem by recognizing that steady states for such models are fixed points of a well-defined equilibrium mapping. This insight allows us to appeal to a relatively newer literature of powerful order-theoretic techniques, tailored for dealing with the monotonicity of fixed points for non-lattice spaces (Acemoglu and Jensen (2015)). A technical challenge that lies with misspecified MDPs is that the underlying spaces containing such fixed points are typically not lattices in any natural order, and therefore, a wide array of popular lattice-dependent techniques (Hopenhayn and Prescott (1992); Topkis (1998)) are not applicable. Finally, our results for comparative statics do not necessitate any specific knowledge of the functional forms of any environment primitives, and therefore are robust, and hold for a very general class of dynamic programming environments with misspecified models. [^7] The reader may be interested in the following discussion: https://stats.stackexchange.com/questions/274815/why-should-i-be-bayesian-when-my-model-is-wrong. Uppal and Wang (2003) discuss the contrast between Savagian and Knightian approaches to tackle model misspecification concerns. The interested reader is referred to their p. 2469 for further reading.
@@
To interpret our framing, we discuss examples that cover three important economic environments. The first example concerns Bayesian inference with misspecified AR(1) models. The agent’s set of models is misspecified in the sense that although the true process is AR(1), the innovations of the process are non-normally distributed and, therefore, are not contained in their set of AR(1) models with Gaussian innovations. We characterize the Berk-Nash equilibrium for this pathological example, and show that despite being misspecified, the Bayesian agent correctly infers the persistence parameter of the true process at the steady state. The next two examples are based on EP, albeit with slight modifications for interpretation and analytical tractability. In the second example, we study a canonical problem of effort provision in a dynamic effort-task problem where the agent seeks to learn their ability while exercising control over the stochastic outcome of a task by choosing their effort levels. The agent’s models are misspecified since they incorrectly postulate that tomorrow’s outcome depends only on their ability and not on today’s outcome. In contrary, the true dynamics are determined by both ability and current outcome. The Berk-Nash equilibrium for this case is the steady-state frequency of the outcome, either a success or a failure, along with the agent’s inference of their ability. We characterize the dependence of these equilibrium objects on the economic primitives of the environment, such as the cost of exerting higher effort or the objective probability of success, and study their associated comparative statics. The final example considers an agent that optimally chooses its consumption and savings in the presence of preference and productivity shocks. The agent seeks to learn about the return of their wealth process but does so through a set of misspecified models that do not consider the correlation between the two shocks. Here, the Berk-Nash equilibrium is the long-run perceived distribution of the wealth process along with their corresponding inferred parameter of the return on that process. However, this case has no analytical solution; therefore, making our results most valuable for such settings.
@@
We now turn to the specificity of our results. Theorem 1 establishes the existence of a Berk-Nash equilibrium for a misspecified MDP with compact action and state spaces. EP and Anderson, Duanmu, Ghosh, and Khan (2023) (hereafter, ADGK) provide proofs for the existence of a Berk-Nash equilibrium with finite and infinite state and actions, respectively. ADGK applies recent advances in non-standard analysis to present existence theorems for misspecified environments with infinite action and state spaces and unbounded payoff functions. We adopt the more general environment of ADGK, but the novelty of our existence result lies in exploiting the non-lattice structure of our economic environment. In general, distributions over states and actions are not lattices under most natural orders; therefore, to show the existence of the Berk-Nash equilibrium which in turn is a distribution, we rely on a novel application of techniques tailored for non-lattice spaces. These techniques, originating in Smithson (1971), have been pioneered in Acemoglu and Jensen (2015) for establishing the existence and comparative statics results in infinite-horizon large dynamic economies.
@@
After furnishing the existence result for the Berk-Nash equilibrium, we utilize these techniques for Theorems 2-4 and give sufficient conditions for which the Berk-Nash equilibrium and the corresponding inferred best-fit model respond monotonically to the primitives of the environment. We formalize a notion of a positive shock for a misspecified dynamic optimization problem, defined as shocks to the primitives that increase strategies for any given beliefs over models, and under mild assumptions establish in Theorem 2 that the least and the greatest inferred best-fit model at the steady state, responds monotonically to a change in the economic primitives. Theorem 2 further provides a prediction for the monotonicity of the Berk-Nash equilibrium in the usual stochastic order, when a misspecified agent is hit with positive shocks to their primitives. In Theorem 3, we show that the least and the greatest inferred best-fit model responds monotonically to an increase in the set of models, where an increase is defined in the strong-set order in the space of parameters. Theorem 4 drops the usual stochastic order on the underlying space and extends Theorem 2 to spaces with increasing and convex orders. [^8] As Che, Kim, and Kojima (2021) note in the context of individual choices, the strong set order proves to be an appropriate notion; see the references therein for further details.
@@
Our final result is of a slightly different flavor from the previous results. In Theorem 5, we perform a welfare comparison of steady-state learning in a correctly specified vis-à-vis a misspecified learning environment. While welfare under misspecified learning is always weakly lower than that under correct learning, we provide an upper bound on the cost of misspecification, formalized as the difference between the two instances and outline its dependence in terms of the primitives of the environment. We also note the potential usefulness of this bound for computational applications concerning Berk-Nash equilibria.
@@
Within the dynamic programming framework, following EP and ADGK, this paper adds to this literature a more general framework to model MDPs with misspecified learning for infinite environments. [^9] See Puterman (1994) for illustrations covering operations research, engineering, and many other allied fields. Also, see Rust (1994) for an extensive survey on MDPs and associated structural estimation methods. The results in this paper could potentially be admissible in an extended learning version of the partially observable MDP framework of Saghafian (2018); see the references therein for an overview of that literature. MDPs are commonly used to analyze instances involving agents making decisions sequentially in an uncertain environment and, therefore, have found many applications within natural and social sciences. Within the economics literature, one of its earliest uses is traceable in Arrow, Harris, and Marschak (1951), where an inventory holder maximizes some given objective (profits, revenue, net utility) by choosing an inventory policy subject to stochastic product demand and other random fluctuations to the primitives of the problem. [^10] This work has important precursors in Arrow, Blackwell, and Girshick (1949). Since then, MDPs have spanned both micro and macro environments, covering many important economic applications; these include (i) investment with adjustment costs under uncertain demand (Lucas and Prescott (1971)), (ii) one-sector optimal economic growth with uncertainty (Brock and Mirman (1972)), (iii) equilibrium search and unemployment (Lucas and Prescott (1974)), and (iv) asset prices in a pure exchange economy (Lucas (1978)). [^11] Other prominent examples include models of portfolio choice under uncertainty (Phelps (1962); Levhari and Srinivasan (1969)) and business cycle models (Kydland and Prescott (1982); Long Jr. and Plosser (1983)). In a recent paper, Saghafian (2018) presents a framework for an MDP in which agents possess a ‘cloud’ of models and allows for considering model ambiguity and the potential misspecification within an MDP.
@@
This paper also contributes to the understanding of the comparative static properties of the limit posterior of a Bayesian inference process. In his foundational paper on inference with misspecified models, devoid of any actions, Berk (1966) demonstrates that for a misspecified Bayesian agent learning about a parameter through a series of independently and identically distributed (i.i.d.) signals, the posterior concentrates asymptotically on those set of models (referred therein as the asymptotic carrier) where the Kullback-Liebler divergence is minimal with respect to the true model. [^12] See also Bunke and Milhaud (1998) and Shalizi (2009) for extensions to non-iid environments. The reader should note that Berk’s asymptotic carrier is independent of the prior distribution. In this paper, we address a misspecified MDP, a non-i.i.d. environment, that combines decision-making and inference but heavily relies on Berk’s characterization to abstract away from the dynamics of the updating process. This approach enables us to focus on the comparative statics of the steady-state behavior for both the Berk-Nash equilibrium and the corresponding asymptotic carrier, and contribute to the inference literature by outlining the dependence of the inferred models on the primitives of the decision making and the learning environment. Furthermore, our results also have implications for a series of papers that follow the papers of Huber (1967) and White (1982). These papers employ maximum likelihood techniques for inference and show that when an agent conducts inference using maximum likelihood estimation, their inference converges towards models that minimize the Kullback–Leibler divergence with respect to the true distribution, identical to the limits proposed in Berk (1966). [^13] For Bayesian inference, Grünwald and van Ommen (2017) classifies misspecification into good or bad, depending on the structure of the models. They further show that this has important implications for the consistency of Bayesian inference.
@@
Naturally, this work is informed by the substantive literature on monotone comparative statics. The theory of monotone comparative statics deals with characterizing conditions under which the optimizing behavior of agents leads to solution concepts (and invariant distributions) being monotonic in the primitives of the environment. [^15] Some important references include Milgrom and Shannon (1994), Amir (1996), Topkis (1998), Huggett (2003), Datta, Reffett, and Woźny (2018), and Light (2021). Amir (2018) is an useful guide for some of the history and also of the more recent advances in the comparative statics literature. In an influential paper, Hopenhayn and Prescott (1992) uses the tools of monotone comparative statics to analyze stationary dynamic optimization problems with lattice environments. There results turn out to be of limited use for obtaining our theorems. Our proofs instead exploit the non-lattice structure of our setting and, therefore, depend on the techniques developed in Smithson (1971) and Acemoglu and Jensen (2015). We adapt and apply these techniques in a novel way to misspecified MDPs with Bayesian learning.
@@
Outline. The paper is organized as follows. Section 2 sets up the general framework for a misspecified MDP and outlines the necessary prerequisites and techniques for the comparative statics analysis. Section 3 offers three examples to illustrate our setting. The main results of this paper are presented in Section 4, while Section 5 completes the formal analysis of the examples. Section 6 compares welfare between correctly specified and misspecified settings. Section 7 concludes by discussing how these results can be extended and related to different settings. Appendix A to this paper contains the necessary mathematical preliminaries, and the proofs of the results are presented in Appendix B.

### v2

# This Version: March 14, 2025

@@
This Version: March 14, 2025
We present novel monotone comparative statics results for steady-state behavior in a dynamic optimization environment with misspecified Bayesian learning. Building on Esponda and Pouzo (2021), we analyze a Bayesian learner whose prior is over parameterized transition models but is misspecified in the sense that the true process does not belong to this set. We characterize conditions that ensure monotonicity in the steady-state distribution over states, actions, and inferred models. Additionally, we provide a new monotonicity-based proof of steady-state existence, derive an upper bound on the cost of misspecification, and illustrate the applicability of our results to several environments of general interest.
Journal of Economic Literature Classification Numbers: C61, D83, D90
Key Words: Berk-Nash equilibrium, monotone comparative statics, model misspecification.
In general, dynamic programs with learning are intractable, i.e., they are not solvable either analytically or numerically, when there is no separability of control and learning. The problem is not only whether a solution exists, but if a solution can be characterized and its properties studied. Koulovatianos, Mirman, and Santugini (2009)
@@ 1 Introduction
Stemming from the foundational contributions of Arrow and Green (1973) and Nyarko (1991), model misspecification in economic environments continues to be of substantive interest for economists. It acknowledges the possibility that economic agents, perhaps due to cognitive limitations or simplified perspectives, may often not know the true model of their complex economic environment and, therefore, learn about it through a set of models that are misspecified in the sense that the true model is not contained in their set of models. A growing body of literature in economics suggests that enriching classical economic environments with misspecified economic agents aligns theoretical predictions closely with the observed behavior. [^1] For example, Farmer, Nakamura, and Steinsson (2024) find that much of the anomalies concerned with forecasts of professional forecasters can be attributed to them not knowing the true model of the environment.
Dynamic programming offers one robust framework for economists to approach discrete-time economic problems. They are commonly used to analyze instances where economic agents make decisions sequentially in an uncertain environment. The framework applies to a broad spectrum of economic models, from consumption-savings decisions for individuals to investment choices for firms. However, they are known to be intractable, no closed-form solutions, when incorporated with agents who simultaneously make decisions and learn about their environment through their models, where both the decision making and the learning activity mutually affect each other. In this paper, we provide robust predictions on the monotone comparative statics properties of an agent’s steady-state behavior within a dynamic programming environment, particularly when they are learning with potentially misspecified models. [^2] As is well-known, monotone comparative statics analysis deals with characterizing conditions for various environments, both static and dynamic, under which decision rules and solution concepts are increasing in the primitives.
@@
$V(s,\mu)=\max_{x\in\mathbb{X}}\Bigl{\{}\int_{\mathbb{S}}\{u\left(s,x,s^{\prime}\right)+\beta V\left(s^{\prime},\mu^{\prime}\right)\}{\bar{Q}_{\mu}}\left(ds^{\prime}\mid s,x\right)\Bigl{\}},$ (1)
@@
For a misspecified agent that follows the above setting in choosing their sequence of actions and updating their sequence of beliefs over models, EP predicts its steady-state behavior in terms of a solution concept called the Berk-Nash equilibrium. It is an equilibrium distribution over states and actions corresponding to which the misspecified agent infers the model (or a set of models) that ‘best-fits’ the true model. In turn, given the inferred model, the chosen actions are optimal for each state. Furthermore, the steady-state is stationary in the sense that equilibrium distribution over states and actions leads to a stationary Markov process over the states. The ‘best-fit’ formalization is in the sense of a minimum weighted Kullback-Leibler divergence, with weights determined by the equilibrium distribution over states and actions. [^4] The Kullback-Leibler (KL) divergence measures the difference between two probability distributions in terms of relative entropy; in our case, between the model-implied process and the true process; see Cover and Thomas (2005) for details. Given this prediction of the long-run behavior of a misspecified Bayesian agent in terms of a Berk-Nash equilibrium and its corresponding inferred model, we ask the following comparative statics question: how do the equilibrium objects respond to changes in the economic primitives? Phrased differently, what are the requirements on the primitives of the economic environment so that a misspecified agent’s stationary distribution and corresponding best-fit models exhibit monotone comparative statics properties?
@@
This question is of general theoretical interest for several reasons. First, models of dynamic programming incorporated with learning, whether correctly specified or misspecified, are known to be intractable. [^5] With learning incorporated in dynamic programming models, the agent is an active learner and processes incoming data in order to infer the models of their unknown environment. At the same time, they also take decisions. These two functions of the agent are intertwined. In an important paper on learning in stochastic growth models, Koulovatianos, Mirman, and Santugini (2009) (hereafter, KMS) note that the intractability arises primarily for two reasons: (a) the curse of dimensionality problem, which occurs because of incorporation of beliefs about the models in the state space, thus making the state space very large and computationally cumbersome, and (b) with Bayesian updating, the prior and the posterior over models potentially belong to different families, thereby hampering analytical and computational tractability. [^6] As KMS and EP note, the curse of dimensionality does not arise if the beliefs are over a finite parameter set. However, for parameter spaces such as the real line, it is a cause for concern. Also note that, unlike KMS, in our framework, choosing actions also affects the flow of information, thus further complicating matters. By focusing on the long-run prediction for a misspecified dynamic program with a precise characterization of the distribution over states and actions, as well as the limiting posterior (beliefs) over models, we abstract away from the intractability problems and thereby make predictions about their steady-state properties even when their closed-form solutions may not exist. Thus, in this paper, we expand the theory of dynamic programming with misspecified learning, potentially rendering it applicable across a wide spectrum of economic domains.
@@
Second, comparative statics for misspecified learning environments are typically not well-understood in the existing literature on misspecification in economic theory. This traces itself to the issue that with concurrent decision making and learning, there is often no separability between the two, as is illustrated in Equation (1). We tackle this problem by focusing on the monotone comparative statics properties and recognizing that steady states for such models are fixed points of a well-defined equilibrium mapping. This insight allows us to appeal to a relatively newer literature of powerful order-theoretic techniques, tailored for dealing with the monotonicity of fixed points for non-lattice spaces (Acemoglu and Jensen (2015)). A technical challenge that lies with misspecified MDPs is that the underlying spaces containing such fixed points are typically not lattices in any natural order, and therefore, a wide array of popular lattice-dependent techniques (Hopenhayn and Prescott (1992); Topkis (1998)) are not applicable. Finally, our results for comparative statics do not necessitate any specific knowledge of the functional forms of any environment primitives, and therefore are robust, and hold for a very general class of dynamic programming environments with misspecified models. [^7] The reader may be interested in the following discussion: https://stats.stackexchange.com/questions/274815/why-should-i-be-bayesian-when-my-model-is-wrong. Uppal and Wang (2003) discuss the contrast between Savagean and Knightian approaches to tackle model misspecification concerns. The interested reader is referred to their p. 2469 for further reading.
@@
To interpret our framing, we discuss examples that cover three important economic environments. The first example concerns Bayesian inference with misspecified AR(1) models. The agent’s set of models is misspecified in the sense that although the true process is AR(1), the innovations of the process are non-normally distributed and, therefore, are not contained in their set of AR(1) models with Gaussian innovations. We characterize the Berk-Nash equilibrium for this example, and show that despite being misspecified, the Bayesian agent correctly infers the persistence parameter of the true process at the steady state. The next two examples are based on EP, albeit with slight modifications for interpretation and analytical tractability. In the second example, we study a canonical problem of effort provision in a dynamic effort-task problem where the agent seeks to learn their ability while exercising control over the stochastic outcome of a task by choosing their effort levels. The agent’s models are misspecified since they incorrectly postulate that tomorrow’s outcome depends only on their ability and not on today’s outcome. In contrary, the true dynamics are determined by both ability and current outcome. The Berk-Nash equilibrium for this case is the steady-state frequency of the outcome, either a success or a failure, along with the agent’s inference of their ability at the steady state. We characterize the dependence of these equilibrium objects on the economic primitives of the environment, such as the cost of exerting higher effort or the objective probability of success, and study their associated comparative statics. The final example considers an agent that optimally chooses its consumption and savings in the presence of preference and productivity shocks. The agent seeks to learn about the return of their wealth process but does so through a set of misspecified models that do not consider the correlation between the two shocks. Here, the Berk-Nash equilibrium is the long-run perceived distribution of the wealth process along with their corresponding inferred parameter of the return on that process. However, this case has no analytical solution; therefore, making our results most valuable for such settings. [^8] Additionally, in Remark 2 we show the applicability of our results to a dynamic version of a setting involving an over-confident agent who has a misspecified model of their ability, based on Heidhues, Kőszegi, and Strack (2018).
@@
We now turn to the specificity of our results. Theorem 1 establishes the existence of a Berk-Nash equilibrium for a misspecified MDP with compact Euclidean state and action spaces. EP and Anderson, Duanmu, Ghosh, and Khan (2024) (hereafter, ADGK) provide proofs for the existence of a Berk-Nash equilibrium with finite and infinite state and actions, respectively. ADGK applies recent advances in non-standard analysis to present existence theorems for misspecified environments with infinite action and state spaces (compact-metric) and unbounded payoff functions. In contrast, here we focus on compact Euclidean spaces, where existence also emerges as a by-product of the monotonicity assumptions. The novelty of our existence result lies in exploiting the non-lattice structure of our economic environment. In general, distributions over states and actions are not lattices under most natural orders; therefore, to show the existence of the Berk-Nash equilibrium which in turn is a distribution, we rely on a novel application of techniques tailored for non-lattice spaces. These techniques, originating in Smithson (1971), have been pioneered in Acemoglu and Jensen (2015) for establishing the existence and comparative statics results in infinite-horizon large dynamic economies.
@@
After furnishing the existence result for the Berk-Nash equilibrium, we utilize these techniques for Theorems 2-4 and give sufficient conditions for which the Berk-Nash equilibrium and the corresponding inferred best-fit model respond monotonically to the primitives of the environment. We formalize a notion of a positive shock for a misspecified dynamic optimization problem, defined as shocks to the primitives that increase strategies for any given beliefs over models, and under mild assumptions establish in Theorem 2 that the least and the greatest inferred best-fit model at the steady state, responds monotonically to a change in the economic primitives. Theorem 2 further provides a prediction for the monotonicity of the Berk-Nash equilibrium in the usual stochastic order, when a misspecified agent is hit with positive shocks to their primitives. In Theorem 3, we show that the least and the greatest inferred best-fit model responds monotonically to an increase in the set of models, where an increase is defined in the strong-set order in the space of parameterized models. Theorem 4 drops the usual stochastic order on the underlying space and extends Theorem 2 to spaces with increasing and convex orders. [^9] As Che, Kim, and Kojima (2021) note in the context of individual choices, the strong set order proves to be an appropriate notion; see the references therein for further details.
@@
Our final result is of a slightly different flavor from the previous results. In Theorem 5, we perform a welfare comparison of steady-state learning in a correctly specified vis-à-vis a misspecified learning environment. While welfare under misspecified learning is always weakly lower than that under correct learning, we provide an upper bound on the cost of misspecification, formalized as the difference between the two instances and outline its dependence in terms of the primitives of the environment. [^10] We also remark on the potential usefulness of this bound for computational applications concerning Berk-Nash equilibria.
@@
Within the dynamic programming framework, following EP and ADGK, this paper works with a general framework (Euclidean) to model MDPs with misspecified learning for infinite environments. [^11] See Puterman (1994) for illustrations covering operations research, engineering, and many other allied fields. Also, see Rust (1994) for an extensive survey on MDPs and associated structural estimation methods. The results in this paper could potentially be admissible in an extended learning version of the partially observable MDP framework of Saghafian (2018); see the references therein for an overview of that literature. MDPs are commonly used to analyze instances involving agents making decisions sequentially in an uncertain environment and, therefore, have found many applications within natural and social sciences. Within the economics literature, one of its earliest uses is traceable in Arrow, Harris, and Marschak (1951), where an inventory holder maximizes some given objective (profits, revenue, net utility) by choosing an inventory policy subject to stochastic product demand and other random fluctuations to the primitives of the problem. [^12] This work has important precursors in Arrow, Blackwell, and Girshick (1949). Since then, MDPs have spanned both micro and macro environments, covering many important economic applications; these include (i) investment with adjustment costs under uncertain demand (Lucas and Prescott (1971)), (ii) one-sector optimal economic growth with uncertainty (Brock and Mirman (1972)), (iii) equilibrium search and unemployment (Lucas and Prescott (1974)), and (iv) asset prices in a pure exchange economy (Lucas (1978)). [^13] Other prominent examples include models of portfolio choice under uncertainty (Phelps (1962); Levhari and Srinivasan (1969)) and business cycle models (Kydland and Prescott (1982); Long Jr. and Plosser (1983)). In a recent paper, Saghafian (2018) presents a framework for an MDP in which agents possess a ‘cloud’ of models and allows for considering model ambiguity and the potential misspecification within an MDP.
@@
This paper also contributes to the understanding of the comparative static properties of the limit posterior of a Bayesian inference process, specifically a Markov process. In his foundational paper on inference with misspecified models, devoid of any actions, Berk (1966) demonstrates that for a misspecified Bayesian agent learning about a parameter through a series of independently and identically distributed (i.i.d.) signals, the posterior concentrates asymptotically on those set of models (referred therein as the asymptotic carrier) where the Kullback-Leibler divergence is minimal with respect to the true model. [^14] See also Bunke and Milhaud (1998) and Shalizi (2009) for extensions to non-iid environments. The reader should note that Berk’s asymptotic carrier is independent of the prior distribution. In this paper, we address a misspecified MDP, a non-i.i.d. environment, that combines decision-making and inference but heavily relies on Berk’s characterization to abstract away from the dynamics of the updating process. This approach enables us to focus on the comparative statics of the steady-state behavior for both the Berk-Nash equilibrium and the corresponding asymptotic carrier, and contribute to the inference literature by outlining the dependence of the inferred models on the primitives of the decision making and the learning environment. [^15] This also has comparative statics implications for a series of papers that follow the papers of Huber (1967) and White (1982). These papers use maximum likelihood techniques for inference and show that an agent’s inference converges to models minimizing the Kullback–Leibler divergence from the true distribution, consistent with the limits proposed in Berk (1966).
@@
Naturally, this work is informed by the substantive literature on monotone comparative statics. The theory of monotone comparative statics deals with characterizing conditions under which the optimizing behavior of agents leads to solution concepts (and invariant distributions) being monotonic in the primitives of the environment. [^17] Some important references include Milgrom and Shannon (1994), Amir (1996), Topkis (1998), Huggett (2003), Datta, Reffett, and Woźny (2018), and Light (2021). Amir (2018) is a useful guide for some of the history and also of the more recent advances in the comparative statics literature. In an influential paper, Hopenhayn and Prescott (1992) use the tools of monotone comparative statics to analyze stationary dynamic optimization problems with lattice environments. Their results turn out to be of limited use for obtaining our theorems. Our proofs instead exploit the non-lattice structure of our setting and, therefore, depend on the techniques developed in Smithson (1971) and Acemoglu and Jensen (2015). We adapt and apply these techniques in a novel way to misspecified MDPs with Bayesian learning.
@@
Outline. The paper is organized as follows. Section 2 sets up the framework for a misspecified MDP and outlines the necessary prerequisites and techniques for the comparative statics analysis. Section 3 offers three examples to illustrate our setting. The main results of this paper are presented in Section 4, while Section 5 completes the formal analysis of the examples. Section 6 compares welfare between correctly specified and misspecified settings. Section 7 concludes by discussing how these results can be extended and related to different settings. Appendix A to this paper contains the necessary mathematical preliminaries, and the proofs of the results are presented in Appendix B.